OpenStudy (littlebird):

Use De Moivre's theorem to find the power of (2-2i)^5. Answer needs to be in standard form.

OpenStudy (psymon):

Well, we need to turn the original portion, the 2-2i part, into polar form first. Do you know how to do that?

OpenStudy (littlebird):

No

OpenStudy (psymon):

Alright, we need to know a couple of conversions then. So the form we want is this: \[r(\cos \theta + isin \theta)\]This means we need conversions togo froom rectangular numbers to radius and angle. So these are the two conversions we need: \[r=\sqrt{a ^{2}+b ^{2}} \]which you may recognize as a form of pythagorean theorem. The second we need is: \[\tan \theta =\frac{ b }{ a }\] In this problem, a is 2 and b is -2. So using a = 2 and b = -2, do you think you could get me r to start out with?

OpenStudy (littlebird):

r=sqrt(8) tan=-1?

OpenStudy (psymon):

Right, square root 8, which we can simplify to 2sqrt2. And we have tan(theta) = -1. We still need to solve for theta.

OpenStudy (littlebird):

-.785

OpenStudy (psymon):

Well, we want to try to avoid decimals as much as we can. It's usually a good habit to keep numbers as regular degree or radian measures. Do you know what that is in regular radian measures?

OpenStudy (littlebird):

I don't remember how to find it.

OpenStudy (psymon):

Well, it requires some unit circle knowledge. For all trig stuff its best to know the unit circle or have a chart handy. So for us the radian measure we need is 7pi/4. This means we have this for our polar form: \[2\sqrt{2}(\cos \frac{ 7\pi }{ 4}+isin \frac{ 7\pi }{ 4 })\]Do you understand how we got that?

OpenStudy (littlebird):

This is my understanding. That whole formula is in polar form. 2sqrt(2) is r. I found a chart and I think I understand how you found the radians now.

OpenStudy (psymon):

That's good :3 Its just remembering the conversions is all. Okay, now that we have this form we can get our 5th power. This is the formula for it. \[r ^{n}(\cos(n \theta)+isin(n \theta))\] n is the power we want. So what this says is we raise r to the 5th power and we multiply our angles by 5. Think you can do that part?

OpenStudy (psymon):

Stay away from decimals best you can, its definitelty best to avoid decimals when possible.

OpenStudy (littlebird):

128sqrt(2)(cos 35pi/4 + isin 35pi/4)

OpenStudy (psymon):

Awesome :D So that would be your answer :P

OpenStudy (psymon):

Unless it wants it in rectangular form, since this is polar form xD

OpenStudy (littlebird):

Thanks, until now, I had barely any idea what de moivres theorem was.

OpenStudy (psymon):

He has like 50 theorems. A lot of the mathematicians who come up with the things we use have a billion theorems xD But yeah, np :3.